An employee who produces x units per hour earns an hourly wage of y = 0.55x + 12 (in dollars). Find the hourly wage for an employee who produces 20 units per hour.
Answer by checkley71(8403)

Y=,055*20+12 OR Y=11+12 OR Y=$33 PER HOUR